# DAOcc: 3D Object Detection Assisted Multi-Sensor Fusion for 3D Occupancy Prediction

### v2

@@ Abstract
Multi-sensor fusion significantly enhances the accuracy and robustness of 3D semantic occupancy prediction, which is crucial for autonomous driving and robotics. However, most existing approaches depend on large image resolutions and complex networks to achieve top performance, hindering their application in practical scenarios. Additionally, most multi-sensor fusion approaches focus on improving fusion features while overlooking the exploration of supervision strategies for these features. To this end, we propose DAOcc, a novel multi-modal occupancy prediction framework that leverages 3D object detection supervision to assist in achieving superior performance, while using a deployment-friendly image feature extraction network and practical input image resolution. Furthermore, we introduce a BEV View Range Extension strategy to mitigate the adverse effects of reduced image resolution. Experimental results show that DAOcc achieves new state-of-the-art performance on the Occ3D-nuScenes and SurroundOcc benchmarks, and surpasses other methods by a significant margin while using only ResNet50 and 256×704 input image resolution. Code will be made available at https://github.com/AlphaPlusTT/DAOcc.
1 Introduction
@@
3D semantic occupancy prediction (occ) is a critical task in autonomous driving [54, 4, 50] and robotic systems [45, 47, 15, 48], where accurately understanding the environment is essential for safe and efficient navigation. Reliable occupancy prediction requires not just accurate spatial data but also a comprehensive understanding of the environment’s context. Achieving this necessitates the integration of data from multiple sensors. LiDAR provides precise 3D spatial information for obstacle detection, while cameras capture visual details like color and texture for a deeper understanding of the scene. By combining these complementary data sources, the accuracy and robustness of occupancy prediction are significantly enhanced.
@@
In existing works on multi-modal [60, 37, 61, 52, 43, 49, 38] or image-based [4, 51, 44, 13, 50, 24, 57, 42, 53] occ tasks, achieving superior performance often involves using extremely high image resolutions and complex image feature extraction networks (see Table 1 and Table 3), such as using 900×1600 resolution input image and ResNet101 [12] equipped with DCN [8, 66]. However, this approach significantly limits the deployment of top-performing occ models on edge devices due to the high computational demands. In comparison to images, point cloud is much sparser. For example, in the training set of the nuScenes [3] dataset, the maximum number of points in a single frame point cloud is only 34,880, which is equivalent to just 2.4% of the number of pixels in a 900×1600 resolution image. Therefore, how to effectively leverage point cloud data within a multi-modal occ framework remains to be further explored.
@@
Moreover, we observe that most works [60, 37, 61, 52, 43, 49] on multi-modal occ primarily focuses on obtaining more effective fusion features, with insufficient exploration into the forms of supervision for these fused features. Although CO-Occ [38] introduces regularization based on implicit volumetric rendering to supervise fused features, it only utilizes distance ground truth from the original point cloud data, failing to fully exploit the geometry and structure information inherent in point cloud features. In contrast, LiDAR-based 3D detectors [59, 7, 34] effectively leverage this information, achieving significantly better performance in 3D object detection tasks compared to image-based 3D detectors [28, 17, 29, 30]. This observation suggests a new research direction: how to better exploit the unique strengths of point cloud data in multi-modal occ tasks.
@@
To fully exploit the geometry and structure information inherent in point cloud features, we augment the baseline model’s unified BEV features with 3D object detection supervision, thereby enhancing the discriminability of the unified BEV features. This renders the unified BEV features more sensitive to object boundaries and enables our method to avoid the need for complex loss function combinations, as depicted in Figure 1. Additionally, given the sparsity of point clouds, we extend the processing range of the point clouds and employ sparse convolutions [55] to mitigate the computational overhead introduced by this extension. We refer to this approach as BVRE (BEV View Range Extension). BVRE provides a larger BEV field of view, offering more contextual information and mitigating the adverse effects of reduced image resolution. It is worth noting that the 3D object detection supervision is only used as an auxiliary branch during training and can be removed during inference for the occ task.
Extensive experiments are conducted to validate the efficiency and effectiveness of the proposed DAOcc. The results show that DAOcc establishes the new state-of-the-art performance on the Occ3D-nuScenes [44] and SurroundOcc [51] benchmarks, while using only ResNet50 [12] and a 256x704 input image resolution. Specifically, on the Occ3D-nuScenes validation set, DAOcc achieves 53.82 mIoU when using the camera mask during training and 48.2 RayIoU without using the camera mask. Additionally, on the SurroundOcc validation set, DAOcc achieves 45.0 IoU and 30.5 mIoU.
@@
We propose DAOcc, a novel multi-modal occupancy prediction framework that leverages 3D object detection to assist in achieving superior performance while using a deployment-friendly image encoder and practical input image resolution.
We introduce a BEV View Range Extension strategy, which provides a larger BEV field of view, offering more contextual information and mitigating the adverse effects of reduced image resolution.
We establish the new state-of-the-art performance on Occ3D-nuScenes dataset and SurroundOcc dataset, while using only ResNet50 and a 256×704 input image resolution.
2 Related Work
3D Occupancy Prediction: 3D occupancy prediction aims to map all occupied voxels in the environment and assign semantic labels, thus providing more fine-grained perception results.
Camera-based 3D occupancy prediction has gained significant attention due to its cost-effectiveness. Early works [4, 16, 62] lack dense surround occupancy supervision, resulting in sparse occupancy predictions. To obtain dense occupancy prediction, OpenOccupancy [49], SurroundOcc [51] and Occ3D [44] developed methods for generating dense occupancy labels and established benchmarks on their respective proposed datasets. It is worth noting that SurroundOcc does not introduce manual annotation. The emergence of these benchmarks further promoted the development of 3D occupancy prediction. PanoOcc [50] adopts voxel query and a coarse-to-fine approach to learn a unified occupancy representation. FlashOcc [57] uses BEV features with 2D convolution to achieve efficient feature extraction and proposes the Channel-to-Heigh transformation to lift the output of BEV to 3D space. SparseOcc [31] leverages a sparse voxel decoder to reconstruct the sparse geometry of a scene and uses sparse queries to predict masks and labels. However, all camera-based methods inevitably need to cope with the challenges of adverse lighting and weather conditions, and many works inevitably lack robustness due to the ill-posed monocular depth estimation problem [14, 21].
@@ 2 Related Work
While camera-based occupancy prediction has demonstrated promising results, multi-modal approaches offer greater reliability and robustness, making them indispensable for practical applications in autonomous driving and robotics. Due to the fact that cameras are susceptible to changing lighting and weather conditions, OccFusion [37] enhances the accuracy and robustness of the occupancy network by integrating features from LiDAR and radar. In concurrent work, OccFusion [61] projects the preprocessed denser and more uniform point cloud onto the image plane to establish a mapping relationship, and performs deformable attention [67] to fuse the corresponding features. Although OccFusion [61] avoids depth estimation, using deformable attention incurs a greater computational burden [11]. Hydra [52] extends FB-BEV [25] to the camera-radar fusion domain and improves the accuracy of depth estimation with the help of radar features. Co-Occ [38] employs K-nearest neighbor search in selecting neighboring camera features to enhance the corresponding LiDAR features and proposes a regularization based on implicit volume rendering. However, this regularization only utilizes the distance ground truth of the point cloud and does not leverage its intrinsic geometry information. EFFOcc [43] proposes an efficient and lightweight multi-modal 3D occupancy network, but it significantly relies on the 3D object detection pre-training of the LiDAR branch for optimal results. This dependence, however, limits the flexibility of the network structure design. Specifically, to maximize the benefits of the 3D detection pre-training, EFFOcc has to adjust its network architecture to align more closely with well-established LiDAR detection networks. In this work, we introduce a simple yet efficient multi-modal occupancy prediction network, eliminating the need for complex deformable attention [23] as well as image depth estimation during feature fusion, and does not require 3D object detection pre-training. Furthermore, to fully exploit the inherent geometry information in point cloud features, we add supervision of 3D object detection as an auxiliary supervision on the fused features.
@@
Multi-Modal 3D Object Detection: Recent multi-modal 3D object detection methods [33, 19, 20, 6, 46] mainly focus on learning effective BEV feature representations. TransFusion [1] proposes a two-stage transformer-decoder based detection head and applies cross attention to obtain image features for each object query. BEVFusion [33] proposes an efficient and generic multi-task multi-sensor fusion framework which unifies multi-modal features in the shared BEV representation space, and introduces a specialized kernel to speed up the BEV pooling operation. In concurrent work, BEVFusion [26] decomposes the LiDAR-camera fusion into two streams that can independently output perception results, and performs feature fusion after the two streams. DAL [14] follows the concept of ‘Detecting As Labeling’ and decouples the fused features during classification and regression. Specifically, it employs the fused features for classification, whereas it utilizes the point cloud features exclusively for regression. In this work, we incorporate the simple yet effective feature fusion approach of BEVFusion [33] and introduce 3D object detection as an auxiliary branch during training.
3 Proposed Method
3.1 Overall Framework
@@
Our purpose is to fully utilize the inherent geometry and structure information in point cloud features within a multi-modal occupancy prediction framework. Previous multi-modal methods have not fully exploited this and can only achieve superior performance through a more complex image encoder and a larger input image resolution. The overall framework of our proposed DAOcc is illustrated in Figure 2. DAOcc takes surround images and their corresponding time-synchronized point cloud as input, and obtains the features of the image and point cloud through the Camera Encoder and LiDAR Encoder, respectively. The 2D image features are transformed into the 3D voxel space by projection and interpolation. Subsequently, the image and point cloud features in the 3D space are compressed along the height dimension to generate the corresponding BEV(Bird’s Eye View) features. A simple 2D convolution is then applied for feature fusion, and the fully convolutional BEV Encoder encodes these fused features to obtain the final BEV representations. Finally, the occupancy head uses the Channel-to-Height operation [57] to restore the height of the BEV representations, resulting in final 3D voxel space representations that can be utilized for occupancy prediction. These modules collectively form the basic network architecture of DAOcc, which will be elaborated in Sec. 3.2.
@@
On the foundation of the basic network, we introduce the BVRE strategy (Sec. 3.3) to compensate for the information loss that arises from the reduction in image resolution. This strategy aims to enrich the spatial contextual semantic information by broadening the perceptual range from the BEV perspective. Furthermore, to fully leverage the inherent geometry and structure information within point cloud features, we incorporate 3D object detection as auxiliary supervision (Sec. 3.4). This auxiliary supervision not only enhances the discriminability of the fused features but also leads to a very concise overall training loss (Sec. 3.5) for our proposed framework.
3.2 Basic Network
@@
LiDAR Encoder. The method of embedding raw LiDAR points into 3D voxelized features is consistent with SECOND [55]. We first voxelize the point cloud, retaining a maximum of 10 points per voxel, resulting in a 3D voxel grid of size .The feature representation for each voxel is obtained by averaging the features of all points within it. Next, we apply 3D sparse convolutions [55] to encode these voxel features, generating spatially compressed LiDAR voxel features , where represents the feature dimensions.
@@
Camera Encoder. For image feature extraction, taking surround images as input, we first use ResNet50 [12] as the backbone to extract multi-scale features, denoted as , where represent features extracted after downsampling. Then, we employ the Feature Pyramid Network (FPN) [27] as the neck to aggregate these multi-scale features. The output feature map has a shape of , where and represent the input resolution of the image, and and indicate the number of channels and the number of surround images, respectively.
@@
Projection and Interpolation. For image-related occupancy prediction, a key step is to transform image features from 2D image plane to 3D volume space. Most existing methods use monocular depth estimation [49, 43, 13, 57] or deformable attention [51, 44, 50, 61]. However, monocular depth estimation is inherently an ill-posed problem [14, 21], while deformable attention imposes a significant computational burden [11]. Given these limitations, we use a simple yet effective projection and sample approach similar to Harley et al. [10]. Concretely, we first predefine a 3D voxel grid with a shape of , where represents the number of voxels along the -axis. The center point of each voxel is then projected onto the image feature plane, and only points that fall within both the image feature plane and the camera’s field of view are retained. Next, the sub-pixel projection positions of the retained points are bilinearly interpolated to generate the image features corresponding to each voxel. For voxels located in the overlapping viewing regions of the surround cameras, we average the image features from the two corresponding cameras to obtain the final feature for each voxel. The output camera’s voxel features can be denoted as .

### v3

@@ Abstract
Multi-sensor fusion significantly enhances the accuracy and robustness of 3D semantic occupancy prediction, which is crucial for autonomous driving and robotics. However, most existing approaches depend on high-resolution images and complex networks to achieve top performance, hindering their deployment in practical scenarios. Moreover, current multi-sensor fusion approaches mainly focus on improving feature fusion while largely neglecting effective supervision strategies for those features. To address these issues, we propose DAOcc, a novel multi-modal occupancy prediction framework that leverages 3D object detection supervision to assist in achieving superior performance, while using a deployment-friendly image backbone and practical input resolution. In addition, we introduce a BEV View Range Extension strategy to mitigate performance degradation caused by lower image resolution. Extensive experiments demonstrate that DAOcc achieves new state-of-the-art results on both the Occ3D-nuScenes and Occ3D-Waymo benchmarks, and outperforms previous state-of-the-art methods by a significant margin using only a ResNet-50 backbone and 256×704 input resolution. With TensorRT optimization, DAOcc reaches 104.9 FPS while maintaining 54.2 mIoU on an NVIDIA RTX 4090 GPU. Code is available at https://github.com/AlphaPlusTT/DAOcc.
I Introduction
@@
3D semantic occupancy prediction (occ) is a critical task in autonomous driving [1, 2, 3] and robotic systems [4, 5, 6], where accurately understanding the environment is essential for safe and efficient navigation. Reliable occupancy prediction requires not just accurate spatial data but also a comprehensive understanding of the environment’s context. Achieving this necessitates the integration of data from multiple sensors. LiDAR provides precise 3D spatial information for obstacle detection, while cameras capture visual details like color and texture for a deeper understanding of the scene. By combining these complementary data sources, the accuracy and robustness of occupancy prediction are significantly enhanced.
@@
In existing works on multi-modal [9, 10, 11, 12, 13] or image-based [2, 14, 15, 16, 3, 17, 18, 19, 20] occ tasks, achieving superior performance often involves using extremely high image resolutions and complex image feature extraction networks (see Table I), such as using 900×1600 resolution input image and ResNet101 [21] equipped with DCN [22, 23]. However, this approach significantly limits the deployment of top-performing occ models on edge devices due to the high computational demands. In comparison to images, the point cloud is much sparser. For example, in the training set of the nuScenes [24] dataset, the maximum number of points in a single point cloud frame is only 34,880, which is equivalent to just 2.4% of the number of pixels in a 900×1600 resolution image. Therefore, how to effectively leverage point cloud data within a multi-modal occ framework remains to be further explored.
@@
Moreover, we observe that most works [9, 10, 11, 12] on multi-modal occ primarily focuses on obtaining more effective fusion features, with insufficient exploration into the forms of supervision for these fused features. Although CO-Occ [13] introduces a regularization based on implicit volumetric rendering to supervise fused features, it only utilizes distance ground truth from the original point cloud data, failing to fully exploit the geometry and structure information inherent in point cloud features. In contrast, LiDAR-based 3D detectors [25, 26, 27] effectively leverage this information, achieving significantly better performance in 3D object detection tasks compared to image-based 3D detectors [28, 29, 30]. This observation suggests a new research direction: how to better exploit the unique strengths of point cloud data in multi-modal occ tasks.
@@
To fully exploit the geometry and structure information inherent in point cloud features, we enhance the baseline model’s unified BEV features by incorporating 3D object detection supervision, thereby improving the discriminability of the unified BEV features. The detection-to-occupancy supervision injects explicit object-localization cues into the occupancy pipeline, sharpening object boundaries and boosting foreground class performance while keeping inference complexity low—since the detection branch is used only during training and removed at inference. It also eliminates the need for complex loss function combinations, as depicted in Figure 2. Additionally, given the sparsity of point clouds, we extend the processing range of the point clouds and employ sparse convolutions [31] to mitigate the computational overhead introduced by this extension. To address the misalignment between the expanded point cloud range and the occupancy label, we apply a combination of coordinate system transformation and feature interpolation to extract fusion features that are spatially aligned with the occupancy label. We refer to this approach as BVRE (BEV View Range Extension). BVRE provides a larger BEV field of view, offering more contextual information and mitigating the adverse effects of reduced image resolution.
Extensive experiments are conducted to validate the efficiency and effectiveness of the proposed DAOcc. The results show that DAOcc establishes the new state-of-the-art performance on the Occ3D-nuScenes [15] and Occ3D-Waymo [15] benchmarks, while using only ResNet50 [21] with a 256×704 input image resolution. Specifically, on the Occ3D-nuScenes validation set, DAOcc achieves a 2.67 mIoU improvement over SGDOCC [7] when trained with camera mask and a 4.7 RayIoU improvement over ALOcc [8] without camera mask. Additionally, on the Occ3D-Waymo validation set, DAOcc demonstrates a 6.08 mIoU boost compared to BEVFormer-Fusion [15]. Furthermore, with TensorRT optimization, DAOcc reaches 104.9 FPS on an NVIDIA RTX 4090 GPU while maintaining 54.2 mIoU on the Occ3D-nuScenes validation set, significantly surpassing the mIoU of previous state-of-the-art methods.
@@
We propose DAOcc, a novel multi-modal occupancy prediction framework that leverages detachable 3D object detection supervision to improve foreground class discrimination and maintain low inference complexity, thus achieving superior performance while using a deployment-friendly image encoder and practical input image resolution.
We introduce a BEV View Range Extension strategy to mitigate the adverse effects of reduced image resolution, and design a flexible feature processing pipeline consisting of coordinate system transformation and feature interpolation, which eliminates the complex manual feature space alignment.
Our proposed method, DAOcc, establishes the new state-of-the-art performance on the Occ3D [15] benchmark, while maintaining competitive inference speed. With TensorRT optimization, DAOcc can achieve 104.9 FPS while significantly surpassing the mIoU of the previous state-of-the-art methods.
II Related Work
II-A 3D Occupancy Prediction
@@ II Related Work
3D occupancy prediction aims to map all occupied voxels in the environment and assign semantic labels, thus providing more fine-grained perception results. While camera-based 3D occupancy prediction [2, 40, 41, 12, 14, 15, 3, 18, 42, 43, 44, 8, 16, 19, 20, 17] has demonstrated promising results, multi-modal approaches offer greater reliability and robustness, making them indispensable for practical applications in autonomous driving and robotics. Due to the fact that cameras are susceptible to changing lighting and weather conditions, OccFusion [10] enhances the accuracy and robustness of the occupancy network by integrating features from LiDAR and radar. In concurrent work, OccFusion [11] projects the preprocessed denser and more uniform point cloud onto the image plane to establish a mapping relationship, and performs deformable attention [45] to fuse the corresponding features. Although OccFusion [11] avoids depth estimation, using deformable attention incurs a greater computational burden [34]. Co-Occ [13] employs K-nearest neighbor search in selecting neighboring camera features to enhance the corresponding LiDAR features and proposes a regularization based on implicit volume rendering. However, this regularization only utilizes the distance ground truth of the point cloud and does not leverage its intrinsic geometry information. SDGOCC [7] leverages both the geometric and semantic information of point clouds to guide the 2D-3D view transformation, and introduces a fusion-to-occupancy-driven active distillation module that selectively transfers multimodal knowledge to image features based on LiDAR-identified regions. Compared to Co-Occ, SDGOCC additionally introduces point cloud semantic information, but its utilization of point cloud geometry information remains limited—relying solely on sparse point cloud based depth supervision for monocular depth estimation, without addressing the inherent ill-posed problem of this task. In this work, we introduce a simple yet efficient multi-modal occupancy prediction network, eliminating the need for complex deformable attention [39] as well as image depth estimation during feature fusion, and avoiding the requirement for 3D object detection pre-training. Furthermore, to fully exploit the inherent geometry information in point cloud features, we incorporate 3D object detection as auxiliary supervision on the fused BEV features.
II-B Multi-Modal 3D Object Detection
@@
Multi-modal 3D object detection has achieved great success, surpassing camera-based [46, 47] and Lidar-based [48] methods by leveraging semantic and geometric information. Recent multi-modal 3D object detection methods [49, 50, 51, 52] mainly focus on learning effective BEV feature representations. TransFusion [53] proposes a two-stage transformer-decoder based detection head and applies cross attention to obtain image features for each object query. BEVFusion [49] proposes an efficient and generic multi-task multi-sensor fusion framework which unifies multi-modal features in the shared BEV representation space, and introduces a specialized kernel to speed up the BEV pooling operation. In concurrent work, BEVFusion [54] decomposes the LiDAR-camera fusion into two streams that can independently output perception results, and performs feature fusion after the two streams. DAL [32] follows the concept of ‘Detecting As Labeling’ and decouples the fused features during classification and regression. Specifically, it employs the fused features for classification, whereas it utilizes the point cloud features exclusively for regression. In this work, we incorporate the simple yet effective feature fusion approach of BEVFusion [49] and introduce 3D object detection as an auxiliary branch during training.
II-C Detection-Occupancy Interaction
Recent work has explored tight couplings between 3D occupancy estimation and 3D object detection, with mutual benefits. OccupancyM3D [55] uses voxel-wise occupancy as auxiliary dense supervision to enhance monocular 3D detection via richer 3D features and geometry priors. SOGDet [56] adds a semantic-occupancy branch that provides dense scene context to multi-view detectors, boosting localization and robustness. Zheng et al. [57] recently proposed an object-centric occupancy completion approach, focusing on high-resolution modeling of foreground objects. DAOcc is complementary to these lines of work, while differing in both supervision direction and practical emphasis.
III Proposed Method
III-A Overall Framework
@@
Our purpose is to fully utilize the inherent geometry and structure information in point cloud features within a multi-modal occupancy prediction framework. Previous multi-modal methods have not fully exploited this and can only achieve superior performance through a more complex image encoder and a larger input image resolution. The overall framework of our proposed DAOcc is illustrated in Figure 3. DAOcc takes surround images and their corresponding time-synchronized point cloud as input, and obtains the features of the image and point cloud through the Camera Encoder and LiDAR Encoder, respectively. The 2D image features are transformed into the 3D voxel space by projection and interpolation. Subsequently, the image and point cloud features in the 3D space are compressed along the height dimension to generate the corresponding BEV(Bird’s Eye View) features. A simple 2D convolution is then applied for feature fusion, and the fully convolutional BEV Encoder encodes these fused features to obtain the final BEV representations. Finally, the occupancy head uses the Channel-to-Height operation [18] to restore the height of the BEV representations, resulting in final 3D voxel space representations that can be utilized for occupancy prediction. These modules collectively form the basic network architecture of DAOcc, which will be elaborated in Sec. III-B.
@@
On the foundation of the basic network, we introduce the BVRE strategy (Sec. III-C) to compensate for the information loss that arises from the reduction in image resolution. This strategy aims to enrich the spatial contextual semantic information by broadening the perceptual range from the BEV perspective. Furthermore, to fully leverage the inherent geometry and structure information within point cloud features, we incorporate 3D object detection as auxiliary supervision (Sec. III-D). This auxiliary supervision not only enhances the discriminability of the fused features but also leads to a very concise overall training loss (Sec. III-E) for our proposed framework.
III-B Basic Network
III-B1 LiDAR Encoder
@@
The method of embedding raw LiDAR points into 3D voxelized features is consistent with SECOND [31]. We first voxelize the point cloud, retaining a maximum of 10 points per voxel, resulting in a 3D voxel grid of size .The feature representation for each voxel is obtained by averaging the features of all points within it. Next, we apply 3D sparse convolutions [31] to encode these voxel features, generating spatially compressed LiDAR voxel features , where represents the feature dimensions.
III-B2 Camera Encoder
@@
For image feature extraction, taking surround images as input, we first use ResNet50 [21] as the backbone to extract multi-scale features, denoted as , where represent features extracted after downsampling. Then, we employ the Feature Pyramid Network (FPN) [58] as the neck to aggregate these multi-scale features. The output feature map has a shape of , where and represent the input resolution of the image, and and indicate the number of channels and the number of surround images, respectively.
III-B3 Projection and Interpolation
@@
For image-related occupancy prediction, a key step is to transform image features from 2D image plane to 3D volume space. Most existing methods use monocular depth estimation [12, 16, 18] or deformable attention [14, 15, 3, 11]. However, monocular depth estimation is inherently an ill-posed problem [32, 33], while deformable attention imposes a significant computational burden [34]. Given these limitations, we use a simple yet effective projection and sample approach similar to Harley et al. [35]. Concretely, we first predefine a 3D voxel grid with a shape of , where represents the number of voxels along the -axis. The center point of each voxel is then projected onto the image feature plane, and only points that fall within both the image feature plane and the camera’s field of view are retained. Next, the sub-pixel projection positions of the retained points are bilinearly interpolated to generate the image features corresponding to each voxel. For voxels located in the overlapping viewing regions of the surround cameras, we randomly select the image feature from one of the two corresponding cameras to obtain the final feature for each voxel. The output camera’s voxel features can be denoted as .
III-B4 BEV Encoder